OpenStudy (anonymous):

What is the derivative of log (base) 8 of x? Thanks for your help.

OpenStudy (anonymous):

\[\frac{d}{dx}\log_b(x)=\frac{1}{x\ln(b)}\]

OpenStudy (anonymous):

Thanks satellite73. You're awesome.

OpenStudy (anonymous):

its like i dont exist. but thats ok. lol.

OpenStudy (anonymous):

Thanks to genius12 also.

OpenStudy (anonymous):

<3

OpenStudy (anonymous):

hey i gave you the medal first right?

OpenStudy (anonymous):

i think so but i also like the d/dx notation better.